# Constraining parity and Lorentz violations in gravity with future ground- and space-based gravitational wave detectors

### v1

@@ II MODIFIED WAVEFORM OF GWS WITH PARITY- AND LORENTZ-VIOLATING EFFECTS
Here a prime represents the derivative for the conformal time $\tau$ and $\mathcal{H}=a^{\prime}/a$. The four parameters, $\bar{\nu}$, $\bar{\mu}$, $\nu_{A}$, and $\mu_{A}$, are introduced to describe deviations from the standard GW propagation in GR, which can arise from specific modified theories of gravity. These parameters capture three distinct classes of effects: (1) Frequency-independent modifications associated with $\bar{\nu}$ and $\bar{\mu}$, which account for changes in the propagation speed of GWs and additional friction effects. (2) Parity-violating effects described by $\nu_{A}$ and $\mu_{A}$, which induce birefringence phenomena, including amplitude and velocity birefringences, in GW propagation. (3) Lorentz-violating effects, also governed by $\bar{\nu}$ and $\bar{\mu}$, which introduce frequency-dependent damping and nonlinear dispersion of GWs. Forecasts for constraints on the frequency-independent and frequency-dependent components of GW friction achievable by upcoming GW detectors are discussed in Refs. Lin:2024pkr ; Zhang:2024rel , while constrains on birefringences Wang:2020cub ; Califano:2023aji and Lorentz-violating dispersion relation have been carried out in Ref. Mirshekari:2011yq . In this work, we focus on the latter two frequency-dependent categories.
@@ II.1 Modified GW waveforms with parity-violating effects
where $\tau_{e}(\tau_{0})$ is the emitted (arrival) time of GWs, $z=1/a(\tau_{e})-1$ is the redshift of GW event, $u=\pi\mathcal{M}f$, $\mathcal{M}$ is the measured chirp mass of the compact binary system, $f$ is the frequency of GW detected by the GW detectors, and the function $\Theta(1+x)=1+x$ for $x\neq-1$ and $\Theta(1+x)=1$ for $x=-1$. In this work, the cosmology parameter we adopt are $\Omega_{m}=0.315$, $\Omega_{\Lambda}=0.685$, and $H_{0}=67.4\;{\rm km}\;{\rm s}^{-1}\;{\rm Mpc}^{-1}$ Planck:2018vyg .

### v2

@@ II MODIFIED WAVEFORM OF GWS WITH PARITY- AND LORENTZ-VIOLATING EFFECTS
Here, a prime represents the derivative for the conformal time $\tau$ and $\mathcal{H}=a^{\prime}/a$. Here, we would like to address some remarks about the validity and the underlined assumptions of Eq. (2.4). This equation is written under the linear approximation, which implies we have ignored all the high-order metric perturbations on the propagation of GWs. We also restrict us to the transverse and traceless modes of GWs for two reasons. Theoretically, the metric perturbations $h_{ij}$ only contains two degenerate traceless and transverse tensor modes in GR, but when the parity- and Lorentz-violating modifications are included, depending on specific types of the parity and Lorentz violations, $h_{ij}$ may contain extra modes, for example, the scalar or vector modes. These extra modes decouple from the two traceless and transverse tensor modes under the linear approximation and thus do not affect the propagation of the two tensorial modes. Observationally, all the GW signals detected by LIGO/Virgo/KAGRA detectors can be well described by the two tensorial polarizations, with no conclusive evidence for scalar or vector modes KAGRA:2021vkt . In addition, the Lorentz- and parity-violating terms included in this equation are assumed to be subdominant compared to the leading-order GR contributions Zhao:2019xmm , which constrains parameter ranges to avoid ghosts or instabilities.
The four parameters, $\bar{\nu}$, $\bar{\mu}$, $\nu_{A}$, and $\mu_{A}$, are introduced to describe deviations from the standard GW propagation in GR, which can arise from specific modified theories of gravity. These parameters capture three distinct classes of effects: (1) Frequency-independent modifications associated with $\bar{\nu}$ and $\bar{\mu}$, which account for changes in the propagation speed of GWs and additional friction effects. (2) Parity-violating effects described by $\nu_{A}$ and $\mu_{A}$, which induce birefringence phenomena, including amplitude and velocity birefringences, in GW propagation. (3) Lorentz-violating effects, also governed by $\bar{\nu}$ and $\bar{\mu}$, which introduce frequency-dependent damping and nonlinear dispersion of GWs. Forecasts for constraints on the frequency-independent and frequency-dependent components of GW friction achievable by upcoming GW detectors are discussed in Refs. Lin:2024pkr ; Zhang:2024rel , while constrains on birefringences Wang:2020cub ; Califano:2023aji and Lorentz-violating dispersion relation have been carried out in Ref. Mirshekari:2011yq . In this work, we focus on the latter two frequency-dependent categories.
@@ II.1 Modified GW waveforms with parity-violating effects
where $\tau_{e}(\tau_{0})$ is the emitted (arrival) time of GWs, $z=1/a(\tau_{e})-1$ is the redshift of GW event, $u=\pi\mathcal{M}f$, $\mathcal{M}$ is the measured chirp mass of the compact binary system, $f$ is the frequency of GW detected by the GW detectors, and the function $\Theta(1+x)=1+x$ for $x\neq-1$ and $\Theta(1+x)=1$ for $x=-1$. In this work, the cosmology parameter we adopt are $\Omega_{m}=0.315$, $\Omega_{\Lambda}=0.685$, and $H_{0}=67.4\;{\rm km}\;{\rm s}^{-1}\;{\rm Mpc}^{-1}$ Planck:2018vyg  [^2] Here we use the Planck cosmological parameters for consistency with previous results presented in LIGOScientific:2019fpa ; LIGOScientific:2021sio ; Gong:2023ffb ; Wang:2025fhw ; Wang:2020cub ; LIGOScientific:2020tif ; Wu:2021ndf . Using different sets of cosmological parameters can only slightly change the results presented in this paper..